OpenStudy (anonymous):

Solve the exponential equation shown below. 12=10^x+3 -9

OpenStudy (anonymous):

\[12=10^{x+3}-9\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, first add the 9 to both sides

OpenStudy (anonymous):

21=10^x+3

OpenStudy (anonymous):

Right, now take \[\log _{10} \] of both sides

OpenStudy (anonymous):

idk how to do that on this kind

OpenStudy (anonymous):

for 21 is it 1.322?

OpenStudy (anonymous):

and -7 for the other side

OpenStudy (anonymous):

\[\log _{10}21=\log _{10}10^{x+3}\]

OpenStudy (anonymous):

The log and exponential on the right cancel out and you're left with \[\log _{10}21=x+3\]

OpenStudy (anonymous):

is it 1.322=-7

OpenStudy (anonymous):

oh! okay i gotcha

OpenStudy (anonymous):

now do i do log of 10 21?